OpenStudy (anonymous):

A car is traveling at a rate of 90 miles per hour. What is the car's rate in feet per second? How many feet will the car travel in 20 seconds? In your computations, use the fact that 1 mile is equal to 5280 feet.Do not round your

OpenStudy (anonymous):

\[\frac{ 90\ miles }{1\ hour }\ \times\ \frac{ 1\ hour }{ 3600\ seconds }\ \times\ \frac{ 5280\ feet }{ 1\ mile }\] multiply through. notice that all units "cancel out" exept for feet and seconds

OpenStudy (mathmale):

...and thus, you'll now have the car's speed in feet per second instead of in miles per hour. Note: don't drive at 90 mph.